OpenStudy (anonymous):

Determine the minimum value of the function f(x)=4^x-8. How would you know? a) 0 b) -8 c) 4 d) no minimum value exists

jimthompson5910 (jim_thompson5910):

Which function is f(x)? \[\Large f(x) = 4^{x} - 8\] \[\Large f(x) = 4^{x - 8}\]

OpenStudy (anonymous):

f(x) = 4x - 8 y=4x-8 replace f(x) with y x=4y-8 switch values and solve for y. 4y=x+8 y=x/4+2

OpenStudy (anonymous):

f(x) is the first one

jimthompson5910 (jim_thompson5910):

focus on just 4^x what happens to the overall expression of 4^x when x is a large negative number (like x = -100) ?

OpenStudy (anonymous):

when x is large, the expression is very small

jimthompson5910 (jim_thompson5910):

a large negative, yes essentially 4^x will approach 0 when x approaches negative infinity so (4^x) - 8 will approach -8 when x approaches negative infinity ------------------------------------------------------- BUT -8 is NOT the minimum because the function output actually never ever gets to -8. It only gets closer and closer. It turns out that there is no minimum because the curve simply approaches this value in the form of a horizontal asymptote.

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):